# Exploring Multidimensional Checkworthiness: Designing AI-assisted Claim Prioritization for Human Fact-checkers

### v1

@@ 3.3. Prototype and Deploy the Design
After completing a high-fidelity wireframe based on the above design specifications, we implemented design features into a functional software prototype. We then conducted pilot user testing to identify potential usability issues that prevent our design from meeting the research goal and to establish specific tasks and protocols for the formal evaluation studies to be conducted with professional fact-checkers. In this section, we describe the user interface (Section 3.3.1) and the technical implementation (Section 3.3.2). See Appendix E for discussion of preliminary findings from pilot tests.
@@ 3.3.2. Technical implementation
Regarding the other three dimensions annotated in Alam et al. (2021)’s dataset, we use the overall “Needs Verification” for our unidimensional baseline. We describe our baseline condition used for experimental study in Section 3.4.2. The two checkworthy dimensions not used were another variant on likely to cause harm and whether the claim merited government attention. Although the dimensions we selected are highlighted as commonly known and important checkworthy factors, as included in Table 1 and also reported by fact-checkers in Liu et al. (2024) and Procter et al. (2023)’s studies. Compared to a full list of Table 1, other factors are omitted due to the lack of available dataset. Additionally, as part of our research limitations, we discuss this in Section 5.3.
@@
The personalized ranking function multiplies each aforementioned model’s predicted probability score by a user-customized weight in range [0,1] and then aggregates the scores (see $S_{l}$ as a linear weighting function in Equation 1). Based on the user assessment of the relative importance of various dimensions, they can directly influence the ranking results by assigning weights to the customized variable for each model, including semantic search, trained classifiers, and LLM classifiers. To further enhance the sensitivity to weight changes and make the system more responsive to user adjustments, we squared the output for each weighted score (see $S_{s}$ in Equation 2). With $S_{l}$, changes in the score primarily reflect variations in probability, offering limited leverage over user weights. In contrast, $S_{s}$ ensures that the rate of change in the ranking score increases alongside the user weight. Additionally, to satisfy that $S_{s}$ changes more rapidly than $S_{l}$ (i.e., $\frac{\partial S_{s}}{\partial W_{i}}>\frac{\partial S_{l}}{\partial W_{i}}$) and given that both $P_{i}$ and $W_{i}$ range from 0 to 1, $W_{i}$ must fall within the range of $\left(\frac{1}{2P_{i}},1\right]$, for $1\geq P_{i}\geq\frac{1}{2}$. This probability range corresponds to scenarios where the model predicts a positive match for the checkworthy factor. Consequently, the squared weighting function increases sensitivity to user-assigned weights for positive predicted claims. However, we acknowledge that this approach also risks disproportionately diminishing the influence of user weights for negative predicted claims, underscoring the need for further refinement.
@@
We implemented our models on the Streamlit server [^7] https://streamlit.io/ and built the front end using Python, AG-Grid JS [^8] https://www.ag-grid.com/, and CSS. We used OpenAI’s gpt-3.5-turbo [^9] https://platform.openai.com as our LLM to create customized classifiers [^10] While more advanced GPT models now exist, the log probability needed for building the ranking function was only available from gpt-3.5-turbo at the time of our implementation..
@@ 3.4. Evaluate via a Mixed-Method Approach
Given our design prototype and study protocol, we proceeded to conduct a mixed-method evaluation with 16 professional fact-checkers. In this section, we describe our formal evaluation protocol, including the study procedure and tasks (Section 3.4.1), participant recruitment (Section 3.4.3), and data collection, and analysis (Section 3.4.4).
@@ 3.4.1. Study procedure and tasks
Our evaluation protocol includes three phases, as described in Figure 3, including 1) task onboarding; 2) a within-subjects experiment; and 3) task reflection.
@@
During the onboarding phase, we first prepared a tutorial video and a checklist to help fact-checkers understand how to use each feature provided in the tool. Before the experiment phase, participants were asked to fill out a pre-screening survey to measure the perceived importance of the four checkworthiness dimentions (described in Section 3.3.2) with a pre-task interview to talk about their current claim prioritization experience.
Next, we scheduled an online meeting with each participant remotely to conduct the experimental study. The within-subjects experiment required participants to perform claim prioritization using two different interfaces: baseline (unidimensional ranking of claims) vs. treatment (multidimensional checkworthy claim ranking). Participants were randomly assigned to two groups (i.e., Block A or B), each using the two interfaces in a different order. Recall that our claims are drawn from Alam et al. (2021)’s COVID-19 dataset (Section 3.3.2). For participant use, the dataset’s test split was further bifurcated into two portions: one used during the familiarization phase and the other used in our actual experiment. We also used different sets of claims in each interface. While this creates a potential difference due to topical effects between baseline vs. treatment conditions, we believe this was justified by the benefit of preventing any learning or familiarization effects that could occur if we had used the same claims in both conditions.
For the experiment, participants were asked to complete a series of claim prioritization tasks using both interfaces. The tasks involved two stages. In the initial exploratory search, participants identified as many interesting claims as possible, first without customized LLM filters and then after creating them. Then, in a constrained selection phase, participants selected three claims they deemed to be the most checkworthy out of all of the claims they had identified in the exploratory phase. They then were asked to complete a post-system survey to evaluate their claim prioritization satisfaction and search experience.
Finally, during the task reflection, participants were first required to assess their subjective use experience via a post-task survey. Unlike the post-system one, this survey aims to assess fact-checker final willingness for tool adoption in claim prioritization. They then asked to recall on their claim selection experience by looking at their final selected claims and describing reasons for selecting them. Additionally, we conducted a semi-constructed interview, asking them to reflect their overall use experience by comparing the baseline and experimental interfaces. A more detailed task description of our claim prioritization tasks is presented in Appendix A. More details about our pre-/post- interview questions, subjective evaluation metrics in the post-system/task survey, and behavioral measurement across the three-stage experimental procedure are documented in Appendix B.
@@ 3.4.2. Baseline condition
In both the experimental and baseline conditions, participants could access features such as query search, user-generated facets powered by the LLM, and personalized weighting options. The key distinction, however, lies in the ranking method. The baseline condition employs a uni-dimensional checkworthiness ranking system trained on the overall “Needs Verification” label from Alam et al. (2021)’s dataset. This single checkworthiness facet, labeled “checkworthy,” is displayed in the left panel of the interface. Participants could adjust its weight using a slider, similar to how they could prioritize other shared factors like query similarity (i.e. topical relevance to the search query) and LLM-generated facets.
@@ 3.4.3. Participant recruitment
We screened global fact-checking websites listed in the Duke Reporters’ Lab and sent recruitment emails to various organizations. To ensure relevant expertise in using claim-monitoring tools, we specifically targeted organizations that mentioned their partnerships with Meta or other technical fact-checking entities, such as Meedan or Full Fact. Fact-checkers in these organizations are more likely to have experience using claim-monitoring tools provided by their partnered tech companies. We recruited 16 participants, including 14 full-time professional journalists who conduct fact-checks and 2 researchers affiliated with the organization with previous experience as fact-checkers (Table 3). All participants took part in our study remotely via Zoom. Each session lasted approximately 1.5 hours. Participants received Amazon Gift Cards or other redeemable options available in their respective countries as compensation.
We intentionally sampled participants who had experience checking social media claims during the recruitment process. We believe these participants helped provide valuable insights as they are more familiar with different search-related tools to look for claims. From the pre-screening survey, all of our participants have used social media monitoring tools such as CrowdTangle, TweetDeck, Trendolizer, Newswhip, or others. Eleven of them also used tools specifically developed for fact-checking, such as the Meta fact-checking tool, Full Fact Alpha, Meedan Check or others. Additionally, eight participants reported they always found claims using either the social media monitoring or the fact-checking tools. Another eight participants said they frequently found claims from these tools. Regarding how many claims they finally checked were from these tools, two participants mentioned almost all claims they checked come from these tools; eight participants reported a very large portion (60% to 90%); and six participants stated a fair amount (40% to 60%).
@@ 4.1.1. Perceived importance
Measures Four dimensions of checkworthiness
Verifiable M(SD)Median Likely false M(SD)Median Likely harmful M(SD)Median Interest to public M(SD)Median Friedman $X^{2}$
@@
Perceived importance 4.44(1.21) — 5 4.63(0.62) — 5 4.81(0.40) — 5 4.50 (0.82) — 5 1.824(0.609)
Ease of finding 4.13(0.96) — 4 3.69(1.08) — 4 4.31(0.87) — 4.5 4.25(1.00) — 4.5 11.735(0.008)
Criterion accuracy 4.13(1.06) — 4 3.87(0.74) — 4 4.20(0.77) — 4 4.07(0.70) — 4 1.878(0.598)
@@
Weight at selection 0.79(0.27) — 0.93 0.75(0.28) — 0.82 0.64(0.36) — 0.69 0.45(0.41) — 0.29 2.375(0.304)
Overall weight 0.77(0.26) — 0.83 0.73(0.28) — 0.81 0.62(0.31) — 0.69 0.39(0.35) — 0.26 0.980(0.046)
Use frequency 2.69(1.96) — 2.50 2.25(1.24) — 2.00 3.06(3.49) — 1.00 2.13(2.75) — 1.50 6.156(0.612)
@@
4.1.2. Ease of finding and Criterion accuracy
Recall that Ease of finding gets at initial impressions in assessing a checkworthy dimension whereas Criterion accuracy focuses on final assessments post-check (Section 3.4.4). Participants generally agreed that “Verifiable” and “Likely false” claims were more difficult to initially identify ($M_{verifiable}=4.13$, $Mdn_{verifiable}=4$, $M_{false}=3.69$, $Mdn_{false}=4$). Significant differences were found in the Friedman test ($X^{2}=11.735$, $p<0.05$). From the post-hoc test (see Table 11 in Appendix H), “Likely false” was rated significantly lower than “Likely harmful” and “Interest to the public” but had no significant difference from “Verifiable.” Regarding Criterion accuracy, “Likely false” also received the lowest rating ($M=3.89$), but no significant difference was observed ($X^{2}=1.878$, $p>0.05$). These results suggest that among the four dimensions, “Likely false” claims may be the most challenging to find, not only initially, but that initial impressions of likely false claims may also prove incorrect after conducting the fact-check.
@@ 4.1.6. Comparison and conclusions
Self-assessments did not show a high variation across participant perceptions of importance of the four dimensions. However, user interaction logs indicated that “Verifiable” and “Likely harmful” were considered relatively more important as these two dimensions either received the highest average weight or were used most frequently. In contrast, “Interest to the public” was not deemed important or used often, as evidenced by comparing participant self-assessment with their actual behaviors. These results confirm that our participants have different priorities over the checkworthy dimensions.
@@ 4.1.7. Why do fact-checkers have different priorities?
One reason is that fact-checking organizations have different priorities. For example, some participants (P4, 5, and 13) mentioned that their fact-checking organizations primarily check political claims. Another important reason reported is the changing news environment: as news events develop, priorities over checkworthiness dimensions also evolves. Such dynamism is further discussed in Section 5.1.3.
@@ 4.2. Fact-checker Hierarchical Approach for Claim Prioritization
How did participants apply different priorities among the four dimensions in claim selection (RQ1.2)? Sehat et al. (2024) reported the absence of any systematic approach to claim prioritization, based on their interviews with fact-checkers, particularly in prioritizing harmful claims. In contrast, our participants appeared to develop an inherent hierarchical approach to filtering and selecting claims according to the relative importance of different dimensions. This difference likely stems from our study’s inclusion of observation beyond participant self-reporting. In particular, our understanding arose from two distinct sources of evidence: 1) qualitative responses describing how participants believed they would filter and select claims using the four checkworthy dimensions, and 2) observational data showing how participants actually selected claims, through an iterative process that involved applying different checkworthy filters.
@@
Our analysis suggests a three-level hierarchy (Figure 4). First, claim assessment begins with filtering based on prerequisite dimensions: mandatory criteria that must be met. Next, participants assessed other criteria we refer to as important but more contextual. Finally, fact-checkers may also consider other, less important criteria as time allows. The most important dimensions (and time) are thus prioritized while remaining open to other dimensions when possible. To support this finding, we first present qualitative responses and quantitative data detailing this hierarchical process. Following this, we then provide design implications for developing more sophisticated tools to facilitate this process.
@@
Only a few participants (P3, 6, 14) mentioned that they would consider “Likely false” after first filtering “Verifiable” claims, due to their organization partnerships with social media platforms, which prioritize addressing false claims.
@@
To visualize our log data and the weighting patterns described above, we map each participant’s actions onto a step-series diagram over the initial, exploratory claim-finding task. Each diagram shows that participant’s slider weight changes for each of the four checkworthy dimensions natively supported, as well as a fifth checkworthy dimension once they created a custom LLM search filter for it. We show diagrams for all participant in Appendix I but include two illustrative examples below. These examples highlight the hierarchical approach for claim prioritization we observed, and we accompany each diagram with explanations from each participant’s retrospective think-aloud.
@@
Figure 5. Slider weight changes by participant P7 during exploratory search are mapped onto a step-series diagram. The X-axis shows key events during the task. The left Y-axis lists checkworthy dimensions: the four standard dimensions above a fifth LLM-customized dimension created by P7. For each dimension, horizontal bars show slider weight over time for that dimension, where bar opacity varies from transparent to opaque as slider weight varies in range [0,1]. Slider weights are initialized to 0.10 and reset to this value when a customized LLM filter is created. The right Y-axis shows our manual analysis of the hierarchical importance of each dimension (Section 4.2), where P denotes a prerequisite checkworthy dimension, C denotes important but contextual, and L denotes less important.

### v2

@@ 3.3. Prototype and Deploy the Design
After completing a high-fidelity wireframe based on the above design specifications, we implemented design features into a functional software prototype. We then conducted pilot user testing to identify potential usability issues that prevent our design from meeting the research goal and to establish specific tasks and protocols for the formal evaluation studies to be conducted with professional fact-checkers. In this section, we describe the user interface (Section 3.3.1) and the technical implementation (Section 3.3.2). See Appendix E for a discussion of preliminary findings from pilot tests.
@@ 3.3.2. Technical implementation
Regarding the other three dimensions annotated in Alam et al. (2021)’s dataset, we used the overall “Needs Verification” for our unidimensional baseline. We describe our baseline condition used for experimental study in Section 3.4.2. The two checkworthy dimensions not used were another variant on likely to cause harm and whether the claim merited government attention. Although the dimensions we selected are highlighted as commonly known and important checkworthy factors, as included in Table 1 and also reported by fact-checkers in Liu et al. (2024) and Procter et al. (2023)’s studies. Compared to a full list of Table 1, other factors are omitted due to the lack of available datasets. Additionally, as part of our research limitations, we discuss this in Section 5.3.
@@
The personalized ranking function multiplies each aforementioned model’s predicted probability score by a user-customized weight in the range [0,1] and then aggregates the scores (see $S_{l}$ as a linear weighting function in Equation 1). Based on the user assessment of the relative importance of various dimensions, they can directly influence the ranking results by assigning weights to the customized variable for each model, including semantic search, trained classifiers, and LLM classifiers. To further enhance the sensitivity to weight changes and make the system more responsive to user adjustments, we squared the output for each weighted score (see $S_{s}$ in Equation 2).
With $S_{l}$, changes in the score primarily reflect variations in AI-predicted probability, offering limited leverage over user weights. In contrast, $S_{s}$ ensures that the rate of change in the ranking score increases alongside the user weight. Additionally, to satisfy that $S_{s}$ changes more rapidly than $S_{l}$ (i.e., $\frac{\partial S_{s}}{\partial W_{i}}>\frac{\partial S_{l}}{\partial W_{i}}$) and given that both $P_{i}$ and $W_{i}$ range from 0 to 1, $W_{i}$ must fall within the range of $\left(\frac{1}{2P_{i}},1\right]$, for $1\geq P_{i}\geq\frac{1}{2}$. This probability range corresponds to scenarios where the model predicts a positive match for the checkworthy factor. Consequently, the squared weighting function increases sensitivity to user-assigned weights for positive predicted claims (i.e., AI predicts that claims satisfy this criterion). However, we acknowledge that this approach also risks disproportionately diminishing the influence of user weights for negative predicted claims, underscoring the need for further refinement.
@@
We implemented our models on the Streamlit server and built the front-end [^8] Streamlit: https://streamlit.io/, AG-Grid: https://www.ag-grid.com/ using Python, AG-Grid JS, and CSS. We used OpenAI’s gpt-3.5-turbo as our LLM to create customized classifiers [^9] While more advanced GPT models now exist, the log probability needed for building the ranking function was only available from gpt-3.5-turbo at the time of our implementation..
@@ 3.4. Evaluate via a Mixed-Method Approach
Given our design prototype and study protocol, we proceeded to conduct a mixed-method evaluation with 16 professional fact-checkers. In this section, we describe our formal evaluation protocol, including the study procedure and tasks (Section 3.4.1), participant recruitment (Section 3.4.3), baseline condition (Section 3.4.2), and data collection and analysis (Section 3.4.4).
@@ 3.4.1. Study procedure and tasks
Our evaluation protocol includes three phases, as described in Figure 3, including 1) task onboarding, 2) a within-subjects experiment, and 3) task reflection.
@@
During the onboarding phase, we first prepared a tutorial video and a checklist to help fact-checkers understand how to use each feature provided in the tool. Before the experiment phase, participants were asked to fill out a pre-screening survey to measure the perceived importance of the four checkworthiness dimensions (described in Section 3.3.2) with a pre-task interview to talk about their current claim prioritization experience.
Next, we scheduled an online meeting with each participant remotely to conduct the experimental study. The within-subjects experiment required participants to perform claim prioritization using two different interfaces: baseline (unidimensional claim) vs. treatment (multidimensional checkworthy claim ranking). Participants were randomly assigned to two groups (i.e., Block A or B), each using the two interfaces in a different order. Recall that our claims are drawn from Alam et al. (2021)’s COVID-19 dataset (Section 3.3.2). For participant use, the dataset’s test split was further bifurcated into two portions: one used during the familiarization phase and the other used in our actual experiment. We also used different sets of claims in each interface. While this creates a potential difference due to topical effects between baseline vs. treatment conditions, we believe this was justified by the benefit of preventing any learning or familiarization effects that could occur if we had used the same claims in both conditions.
For the experiment, participants were asked to complete a series of claim prioritization tasks using both interfaces. The tasks involved two stages. In the initial exploratory search, participants identified as many interesting claims as possible, first without customized LLM filters and then after creating them. Then, in a constrained selection phase, participants selected three claims they deemed to be the most checkworthy out of all of the claims they had identified in the exploratory phase. They were then asked to complete a post-system survey to evaluate their claim prioritization satisfaction and search experience.
Finally, during the task reflection, participants were first required to assess their subjective use experience via a post-task survey. Unlike the post-system one, this survey aims to assess fact-checker final willingness for tool adoption in claim prioritization. They were then asked to recall their claim selection experience by looking at their final selected claims and describing reasons for selecting them. Additionally, we conducted a semi-constructed interview, asking them to reflect on their overall use experience by comparing the baseline and experimental interfaces. A more detailed task description of our claim prioritization tasks is presented in Appendix A. More details about our pre-/post- interview questions, subjective evaluation metrics in the post-system/task survey, and behavioral measurement across the three-stage experimental procedure are documented in Appendix B.
@@ 3.4.2. Baseline condition
In both the experimental and baseline conditions, participants could access features such as query search, user-generated facets powered by the LLM, and personalized weighting options. The key distinction, however, lies in the ranking method. The baseline condition employs a uni-dimensional checkworthiness ranking system trained on the overall “Needs Verification” label from Alam et al. (2021)’s dataset. This single checkworthiness facet, labeled “checkworthy,” is displayed in the left panel of the interface. Participants could adjust its weight using a slider, similar to how they could prioritize other shared factors like query similarity (i.e., topical relevance to the search query) and LLM-generated facets.
@@ 3.4.3. Participant recruitment
We screened global fact-checking websites listed in the Duke Reporters’ Lab [^10] Duke Reporters’ Lab: https://reporterslab.org/locations/ and sent recruitment emails to various organizations. To ensure relevant expertise in using claim-monitoring tools, we specifically targeted organizations that mentioned their partnerships with Meta or other technical fact-checking entities, such as Meedan or Full Fact. Fact-checkers in these organizations are more likely to have experience using claim-monitoring tools provided by their partnered tech companies. We recruited 16 participants, including 14 full-time professional journalists who conduct fact-checks and 2 researchers affiliated with the organization with previous experience as fact-checkers (Table 3). All participants took part in our study remotely via Zoom. Each session lasted approximately 1.5 hours. Participants received Amazon Gift Cards or other redeemable options available in their respective countries as compensation.
We intentionally sampled participants who had experience checking social media claims during the recruitment process. We believe these participants help provide valuable insights as they are more familiar with different search-related tools to look for claims. From the pre-screening survey, all of our participants have used social media monitoring tools such as CrowdTangle, TweetDeck, Trendolizer, Newswhip, or others. Eleven of them also used tools specifically developed for fact-checking, such as the Meta fact-checking tool, Full Fact Alpha, Meedan Check, or others. Additionally, eight participants reported they always found claims using either social media monitoring or fact-checking tools. Another eight participants said they frequently found claims from these tools. Regarding how many claims they finally checked were from these tools, two participants mentioned almost all claims they checked come from these tools; eight participants reported a very large portion (60% to 90%); and six participants stated a fair amount (40% to 60%).
@@ 4.1.1. Perceived importance
Measures Four dimensions of checkworthiness Friedman $X^{2}$
Verifiable Likely false Likely harmful Interest to public
M(SD) Median M(SD) Median M(SD) Median M(SD) Median
@@
Perceived importance 4.44(1.21) 5 4.63(0.62) 5 4.81(0.40) 5 4.50 (0.82) 5 1.82(0.60)
Ease of finding 4.13(0.96) 4 3.69(1.08) 4 4.31(0.87) 4.5 4.25(1.00) 4.5 11.73(0.00)
Criterion accuracy 4.13(1.06) 4 3.87(0.74) 4 4.20(0.77) 4 4.07(0.70) 4 1.87(0.59)
@@
Weight at selection 0.79(0.27) 0.93 0.75(0.28) 0.82 0.64(0.36) 0.69 0.45(0.41) 0.29 2.37(0.30)
Overall weight 0.77(0.26) 0.83 0.73(0.28) 0.81 0.62(0.31) 0.69 0.39(0.35) 0.26 0.98(0.04)
Use frequency 2.69(1.96) 2.50 2.25(1.24) 2.00 3.06(3.49) 1.00 2.13(2.75) 1.50 6.15(0.61)
@@
4.1.2. Ease of finding
Participants generally agreed that “Verifiable” and “Likely false” claims were more difficult to initially identify ($M_{verifiable}=4.13$, $Mdn_{verifiable}=4$, $M_{false}=3.69$, $Mdn_{false}=4$). Significant differences were found in the Friedman test ($X^{2}=11.735$, $p<0.05$). From the post-hoc test (see Table 11 in Appendix H), “Likely false” was rated significantly lower than “Likely harmful” and “Interest to the public” but had no significant difference from “Verifiable.”
4.1.3. Criterion accuracy
Recall that Ease of finding gets at initial impressions in assessing a checkworthy dimension whereas Criterion accuracy focuses on final assessments post-check (Section 3.4.4). In this final assessment, “Likely false” also received the lowest rating ($M=3.89$), but no significant difference was observed ($X^{2}=1.878$, $p>0.05$). These results suggest that among the four dimensions, “Likely false” claims may be the most challenging to find, not only initially, but that initial impressions of likely false claims may also prove incorrect after conducting the fact-check.
@@ 4.1.7. Comparison and conclusions
Self-assessments did not show a high variation across participant perceptions of the importance of the four dimensions. However, user interaction logs indicated that “Verifiable” and “Likely harmful” were considered relatively more important as these two dimensions either received the highest average weight or were used most frequently. In contrast, “Interest to the public” was not deemed important or used often, as evidenced by comparing participant self-assessment with their actual behaviors. These results confirm that our participants have different priorities over the checkworthy dimensions.
@@ 4.1.8. Why do fact-checkers have different priorities?
One reason is that fact-checking organizations have different priorities. For example, some participants (P4, 5, and 13) mentioned that their fact-checking organizations primarily check political claims. Another important reason reported is the changing news environment: as news events develop, priorities over checkworthiness dimensions also evolve. Such dynamism is further discussed in Section 5.1.3.
@@ 4.2. Fact-checker Hierarchical Approach for Claim Prioritization
How did participants apply different priorities among the four dimensions in claim selection (RQ1.2)? Sehat et al. (2024) reported the absence of any systematic approach to claim prioritization based on their interviews with fact-checkers, particularly in prioritizing harmful claims. In contrast, our participants appeared to develop an inherent hierarchical approach to filtering and selecting claims according to the relative importance of different dimensions. This difference likely stems from our study’s inclusion of observation beyond participant self-reporting. In particular, our understanding arose from two distinct sources of evidence: 1) qualitative responses describing how participants believed they would filter and select claims using the four checkworthy dimensions, and 2) observational data showing how participants actually selected claims through an iterative process that involved applying different checkworthy filters.
@@
Our analysis suggests a three-level hierarchy (Figure 4). First, claim assessment begins with filtering based on prerequisite dimensions: mandatory criteria that must be met. Next, participants assessed other criteria we refer to as important but more contextual. Finally, fact-checkers may also consider other, less important criteria as time allows. The most important dimensions (and time) are thus prioritized while remaining open to other dimensions when possible. To support this finding, we present qualitative responses and quantitative data detailing this hierarchical process.
@@
Only a few participants (P3, 6, 14) mentioned that they would consider “Likely false” after first filtering “Verifiable” claims due to their organizational partnerships with social media platforms, which prioritize addressing false claims.
@@
To visualize our log data and the weighting patterns described above, we map each participant’s actions onto a step-series diagram over the initial exploratory claim-finding task. Each diagram shows that participant slider weight changes for each of the four checkworthy dimensions natively supported, as well as a fifth checkworthy dimension once they created a custom LLM search filter for it. We show diagrams for all participants in Appendix I but include two illustrative examples below. These examples highlight the hierarchical approach for claim prioritization we observed, and we accompany each diagram with explanations from each participant’s retrospective think-aloud.
@@
Figure 5. Slider weight changes by participant P7 during an exploratory search are mapped onto a step-series diagram. The X-axis shows key events during the task. The left Y-axis lists checkworthy dimensions: the four standard dimensions above a fifth LLM-customized dimension created by P7. For each dimension, horizontal bars show slider weight over time for that dimension, where bar opacity varies from transparent to opaque as slider weight varies in the range [0,1]. Slider weights are initialized to 0.10 and reset to this value when a customized LLM filter is created. The right Y-axis shows our manual analysis of the hierarchical importance of each dimension (Section 4.2), where P denotes a prerequisite checkworthy dimension, C denotes important but contextual, and L denotes less important.